OpenStudy (anonymous):

Find the surface area of a right circular cone whose height is 12m and radius is 15m.

OpenStudy (anonymous):

Use the formula 2pi(radius)(height)+2pi(r)^2

OpenStudy (anonymous):

surface area = πr^2 + πrl = πr[r + l] for finding l use the formula l = sqrt(h^2 + r^2) so l = sqrt(12^2 + 15^2) = sqrt(144 + 225) = sqrt(369) = 19.21 using π = 3.14 surface area = 3.14*15m [15m + 19.21m] = 3.14*15m*34.21m = 1611.29 m^2